# Direct Detection of Dark Photon Dark Matter with the James Webb Space Telescope

### v2

@@ 3 The Optical Telescope Element (OTE) of JWST
$V_{1}$(mm) 0 7169.0 -796.3 1047.8
$V_{2}$(mm) 0 0 0 0
$V_{3}$(mm) 0 0 -0.19 -2.36
Size(mm) 6605.2(diameter) 738(diameter) 728(length) $\times$517(width) 172.5(diameter)
Table 1: Parameters for the optical elements of JWST. The data can be found in JWST documentation [61].
In the table, ‘RoC’ denotes the radius of curvature, and ‘conic’ denotes the conic constant $K$ which can be related to eccentricity $e_{\mathrm{cone}}$ as
@@
Figure 1: The OTE (Optical Telescope Element) of JWST. Here ISIM stands for ’Integrated Science Instrument Module’. Everything is drawn to scale. Figure taken from Ref. [56].
@@
Table 1 provides insights into the optical characteristics of the JWST’s optical elements. The primary and tertiary mirrors exhibit elliptical shapes, while the secondary mirror is hyperbolic and the fine steering mirror is flat. The primary, secondary and fine steering mirrors are rounded, while the tertiary mirror is rectangular[62]. Their sizes are listed in Table 1. $V_{1}$, $V_{2}$ and $V_{3}$ are the spacial displacements of the mirrors, as shown in Fig. 1.
The EM energy flux density induced by DPDM originates from several sources within the optical system. DPDM can interact with different components of the optical train, including the primary mirror. The light emitted from the primary mirror is then sequentially reflected by the secondary mirror, the tertiary mirror, and the fine steering mirror before reaching the detector. Additionally, DPDM can also directly interact with the secondary mirror, leading to a ‘double’ reflection before being detected. The same principle applies to interactions with the tertiary and fine steering mirrors. However, it is important to note that not all the lights induced by DPDM on the mirrors will reach the detector. Some of them may stray and cannot be focused on the detector after multiple reflections between the mirrors. Even if the light can reach the detector, it wouldn’t have contribution to the data if it’s direction is different from the reflected star light. After all, the JWST mirror system is designed to focus parallel lights from distant sources, while our light rays are always perpendicular to the mirror surface at which they are induced.
Therefore, in the space mission the light generated on the reflectors can be neglected, however the conversion from DPDM to EM signal also takes place in the interior of the detectors of JWST. In the following section we will use this part of the signal to set constraint on the kinetic mixing parameter $\epsilon$. We also propose to do a ground test, where we can change the relative positions of the mirrors to focus the DPDM induced signal. See section 5 for more details on the ground test with modified JWST configurations.
4 Constraints from JWST Observation Data
In this section, we use JWST observation data to set constraint on the kinetic mixing parameter $\epsilon$. DPDM converts to EM signal on every conductor surface, which include all the reflectors and the interior of the detector. The electromagnetic signal generated by DPDM on the reflector propagates along the normal direction of the reflector, which differs from the path of reflected light. Because JWST is designed to reject stray light outside its precise optical path, signals from DPDM that are not aligned with the optical axis cannot be focused onto the detector and thus remain undetected.
On the other hand, DPDM can also convert to EM signal in the interior of the detector. This part of the signal has contribution to the observed data, since the conversion happens very close to the light-sensing pixels. We estimate the total flux by multiplying the converted flux density $\frac{2}{3}\epsilon^{2}\rho_{\mathrm{DM}}$ with the surface area of the detector. To compare it with the sensitivity of JWST, we introduce an equivalent flux density $I_{\mathrm{eqv}}$ incident at the primary mirror, which can be calculated by dividing the total flux by the area of the primary mirror. The detector of MIRI is a square with side length 2.56cm, so the equivalent flux density is
$I_{\mathrm{eqv}}=\frac{2}{3}\epsilon^{2}\rho_{\mathrm{DM}}\times\frac{(2.56\mathrm{cm})^{2}}{\pi(330.26\mathrm{cm})^{2}}=1.27505\times 10^{-5}\epsilon^{2}\rho_{\mathrm{DM}}$ (4.1)
The detector of NIR camera is a square with side length 3.6864cm. The equivalent flux density is
$I_{\mathrm{eqv}}=\frac{2}{3}\epsilon^{2}\rho_{\mathrm{DM}}\times\frac{(3.6864\mathrm{cm})^{2}}{\pi(330.26\mathrm{cm})^{2}}=3.96591\times 10^{-5}\epsilon^{2}\rho_{\mathrm{DM}}$ (4.2)
Figure 2: Model-independent 95% C.L. upper limits on a constant monochromatic signal from JWST data. The first figure corresponds to NIRSpec observation data, and the lower one MIRI. It shows the strongest limit from the all projects at each frequency bin.
The James Webb Space Telescope (JWST) stands as the cutting-edge space telescope, equipped with various detectors and versatile observation modes. In our study, we harnessed data from 972 distinct observation projects to establish constraints on the Dark Photon-Dark Matter coupling constant. Out of these projects, 713 relied on the Near-Infrared Spectrograph (NIRSpec) [63], while 259 made use of the Mid-Infrared Instrument (MIRI) [64]. It is worth noting that the data selected for analysis excludes background subtraction, ensuring its suitability for our research.
The JWST data we collected from the Mikulski Archive for Space Telescopes (MAST) database [65] includes two crucial parameters: the measured spectral flux density, denoted as $\bar{O}_{i}$, and the associated statistical uncertainty, denoted as $\sigma_{\bar{O}_{i}}$. In our effort to establish upper limits on the coupling of Dark Photon with the Standard Model electromagnetic current, denoted as $\epsilon$, we followed the data analysis approach detailed in previous works [66, 57, 46].
To provide a concise overview, we summarize the key aspects of our method here, while reserving more detailed information for Appendix D. Our analysis begins by applying a local polynomial function to model the background surrounding a selected frequency bin, $i_{0}$, while considering neighboring bins. We estimate systematic uncertainties by comparing data deviations to the background fit. Next, we introduce a hypothetical Dark Photon Dark Matter signal with a strength denoted as $S$ at the specific bin, $i_{0}$. This allows us to construct a likelihood function, $L$, that incorporates $S$ into the comparison between data and the background function. Nuisance parameters are introduced to account for the coefficients of the background polynomial function.
Following the statistical method developed in Ref. [66], we compute the ratio, $\lambda_{S}$, between the maximized likelihood under two conditions: first, when only the nuisance parameters are varied to maximize $L$ while keeping $S$ constant, and second, when both the nuisance parameters and $S$ are varied to maximize $L$. The test statistic, $-2\log(\lambda_{S})$, follows a half-$\chi^{2}$ distribution [66]. This analysis allows us to derive the $95\%$ confidence level upper limit, $S_{\mathrm{lim}}$, for a constant monochromatic signal. The results are illustrated in Fig. 2.
We establish upper limits on the mixing parameter $\epsilon$ as $S_{\mathrm{lim}}=S^{\prime}_{\mathrm{eqv}}$, where $S^{\prime}_{\mathrm{eqv}}$ represents the signal strength from theoretical calculations for Dark Photon Dark Matter. Different datasets from NIRSpec and MIRI yield varying constraints on the signal strength coupling $\epsilon$, and we select the most stringent among them. The constraint on $\epsilon$ and its comparison with previous experiments are illustrated by red solid curve in Fig. 4.
@@ 5 Projected Sensitivity with a Modified JWST Configuration
The arrangement of JWST’s mirrors has been optimized for observing distant astrophysical objects, making it not ideal for the purpose of observing DPDM signal. This is the main reason that the constraints obtained using JWST data are relatively weak. Therefore, we propose to use the JWST primary mirror for a ground test. If we can adjust the relative positions of the mirrors, it is possible to focus the DPDM-induced signal generated on the primary mirror onto the detector. We have analyzed this possibility and projected sensitivity limits under the modified mirror configuration. We will demonstrate that the state-of-the-art technology employed in JWST is exceptionally powerful for detecting novel or unconventional signals.
@@
Firstly, we can alter the direction of light rays to make them move to the right for convenience, while at the same time we replace each reflector with a corresponding type of lens as depicted in Figure 3. The radii of curvature for the first three lenses are denoted by $\rho_{1}$, $\rho_{2}$, and $\rho_{3}$, respectively. In this section, we will use subscripts $P$, $S$, $T$ and $F$ for the shorthand of “primary mirror”, “secondary mirror”, “tertiary mirror” and “fine steering mirror” respectively.
In ray-transfer-matrix method, a ray is described with a 2-component vector $X$, the first component being the angle between the ray and the optical axis, the second being its vertical displacement from the optical axis. Each optical operation that a ray undergoes—such as free travel or refraction through a lens—is represented by a $2\times 2$ matrix. Specifically, within this section, free travel over a distance $L_{i}$ will be symbolized by the matrix $U_{i}$, while refraction on the primary mirror will be denoted by the matrix $U_{P}$, and likewise for other mirrors in the optical system.
@@
Light is emitted from each reflector, since the primary mirror is largest in size, we focus on the light generated on the primary mirror. Thex corresponding vector is
@@
Here $y$ represents the height of the emission point relative to the optical axis.
@@
Since the light from distant stars are transformed into a thin beam of light before being picked up by the detector, we should require the same for the DPDM signal. This means, after a series of reflections and propagation, the first component of the ray vector $U_{T}U_{2}U_{S}U_{1}X_{P}$ should be independent of $y$. This gives
@@
We also hope that the signal is not lost in the middle, this means all the light generated on the primary mirror should be picked up by the secondary mirror, tertiary mirror and fine steering mirror. This gives us a set of equations,
@@
Where $D_{P}$, $D_{S}$ and $D_{F}$ is the diameters of the primary mirror, secondary mirror and fine steering mirror, $A_{T}$ and $B_{T}$ are the length and width of the tertiary mirror. Combining Eq. (5.4), we find that these inequalities are satisfied with
@@
When all the signal generated on the whole primary mirror can be detected, it’s straight forward to determine the equivalent flux,
@@
This value, divided by the bandwidth, should be compared with the designed sensitivity of spectral flux density. The projection curve is shown by red dashed curve in Fig. 4. Since the faintest signal in JWST data is from the observation of the cold sky, we directly use this data as an estimation of the future ground test background, assuming the ground test is performed with a background temperature comparable to the cosmic radio background. A lower background temperature would give us a stronger constraint.
Figure 4: Constraints on the kinematic mixing parameter $\epsilon$ between Dark Photon Dark Matter (DPDM) and photons in the randomized polarization scheme. The solid red curve represents the $95\%$ confidence level (C.L.) exclusion limit for DPDM using JWST data. The dashed red curve shows the projected sensitivity for the modified JWST configuration. The left and right sections display constraints derived from NIRSpec and MIRI observation data, respectively. Additionally, we provide a comparison with existing limits, including those from Solar [67], XENON1T [68], Lampost [69], Mudhi [70], Funk [54], and Tokyo [50].
@@ 6 Summary and Outlook
In this study, we conducted a direct detection search for DPDM using a haloscope setup. We explore direct detection of local DPDM through its conversion into a conventional electromagnetic field at JWST’s pixel detectors. Due to the frequency match between the electromagnetic field and the DPDM mass, the resulting signal took the form of a monochromatic electromagnetic wave.
We utilized data from JWST observations to search for a monochromatic signal within the continuous background in the 10-500 terahertz (THz) range. Both the JWST Mid-Infrared Instrument (MIRI) and Near Infrared Spectrograph (NIRSpec) data were employed. Our analysis enabled us to establish limits on the DPDM kinetic mixing coupling at approximately $\epsilon\sim 3\times 10^{-11}$ and $\epsilon\sim\mathcal{O}(10^{-12})$ respectively. This broadband search for DPDM yielded valuable lower-frequency constraints, complementing other experiments conducted in room-sized laboratories, such as Lampost, Mudhi, FUNK, and TOKYO. However, our results indicated a coupling weaker by about one order of magnitude compared to the XENON1T results, which utilized the potential dark photon flux generated by the Sun, without assuming that the dark photon is the dark matter. Our results are also comparable with the astrophysical bound from the solar emission of the dark photon particles.
While our results are slightly weaker than the current XENON1T constraints, JWST shows remarkable potential, especially considering that it was not designed for the direct detection of DPDM. Thanks to its highly sensitive receiver capable of detecting THz signals, JWST provides a strong potential for novel signal searches. By adopting a modified mirror configuration, the sensitivity to DPDM-induced signals can be enhanced by orders of magnitude through proper focusing. Moreover, we demonstrated that for high-mass DPDM, the generated signal reaching the pixel detector follows simple ray optics, greatly simplifying the calculation.

### v3

@@ 3 The Optical System of JWST
$V_{1}$($m$) 0 7169.0 -796.3 1047.8
$V_{2}$($m$) 0 0 0 0
$V_{3}$($m$) 0 0 -0.19 -2.36
Size($m$) 6605.2(diameter) 738(diameter) 728(length) $\times$517(width) 172.5(diameter)
Table 1: Parameters for the Optical Telescope Element (OTE) of JWST. The data can be found in JWST documentation [56].
Figure 1: The OTE (Optical Telescope Element) of JWST. Here, ISIM stands for ’Integrated Science Instrument Module’. Everything is drawn to scale. Figure taken from Ref. [36].
Firstly, let’s focus on OTE. The OTE of JWST comprises a primary mirror, a secondary mirror, a tertiary mirror, and a fine steering mirror. A sketch of the mirror system is shown in Figure 1. Detailed parameters for these optical components are available in JWST documentation [56], and we have also summarized them in Table 1. In the table, ‘RoC’ denotes the radius of curvature, and ‘conic’ denotes the conic constant $K$ which can be related to eccentricity $e_{\mathrm{cone}}$ as
@@
Table 1 provides insights into the optical characteristics of the JWST’s optical elements. The primary and tertiary mirrors exhibit elliptical shapes, while the secondary mirror is hyperbolic, and the fine steering mirror is flat. The primary, secondary, and fine steering mirrors are rounded, while the tertiary mirror is rectangular [55]. Their sizes are listed in Table 1. $V_{1}$, $V_{2}$, and $V_{3}$ are the spatial displacements of the mirrors, as shown in Fig. 1.
As we have argued in the previous sections, dark photons interact with metal surfaces and convert into photon signals. This conversion can occur on the four mirrors in OTE. In the space mission scenario, JWST is optimized to detect light from target stars, meaning that only photons aligned with the reflected starlight path contribute to the detected signal. However, photons produced by DPDM on any mirrors propagate along the local surface normal, which deviates from the starlight optical path. As a consequence, these photons cannot be brought to focus and do not significantly contribute to the observed signal.
Secondly, we focus on the detector optics. Beyond the mirrors in OTE, there are several optical elements inside the detector system where the DPDM-photon conversion could also occur. Consider, for example, the NISpec detector optics shown in Fig. 2. Dark photons may convert into photons on any element along the optical path; however, whenever this occurs on a reflecting mirror, the converted photons do not contribute to the signal because they propagate along the local surface normal rather than the reflected starlight path, and therefore cannot be focused. In addition to reflecting mirrors, the remaining elements include the filter wheel and MSA, IFU+FS. These components are perpendicular to the optical path, so one might expect that DPDM-induced photons on their surfaces could finally contribute to the detected signal. To be more specific, “MSA, IFU+FS” refers to the micro-shutter assembly (MSA) together with the small apertures for the integral field unit (IFU) and fixed slits (FS) [51, 50]. The MSA is of particular interest: it comprises many shutters, each of which can be opened or closed to admit or block light. The shutters are metallic, so DPDM can in principle induce photon signals on them. During observations, only a small fraction of shutters are open while the majority remain closed to isolate selected targets. DPDM could generate photons on these closed shutters; since these shutters on MSA are perpendicular to the light path, one may expect that the photons generated by DPDM will finally contribute to the detected signal. However, this is also impossible, as explained below.
[…]
Thirdly, one might consider the possibility that dark photons could directly induce signals on the detector pixels, thereby bypassing losses along the optical path. However, a careful examination also rules out this contribution. For starlight photons, the light is dispersed by the grating wheel before reaching the detector, such that the signals on different pixels are from different frequencies. This design enhances the spectral resolution. In contrast, dark photons induce photons of the same frequency across all pixels. As a result, the detector responds very differently to DPDM-induced photons on the pixels compared to starlight photons transmitted through the grating. This mismatch effectively eliminates the frequency-resolving power for DPDM-induced photons, significantly weakening their contribution to the signal.
The above analyses indicate that the current configuration of JWST in space is unlikely to detect the DPDM-induced photons on the telescope. However, an alternative approach can be considered: using JWST in a ground-based test setup, where the positioning of its mirrors can be moderately adjusted. This flexibility allows one to focus the light emitted from the primary mirror onto the detector. In this configuration, DPDM-induced photons emitted from the primary mirror would be sequentially reflected by the secondary mirror, tertiary mirror, and fine steering mirror before reaching the detector. While DPDM can also induce signals via direct interactions with the secondary, tertiary, and fine steering mirrors, their contributions would be negligible due to their significantly smaller surface areas compared to the primary mirror. Importantly, this setup requires only adjustments to the mirror positions within the OTE system, while leaving the detector optics untouched. This is because the DPDM-induced photons would behave identically to ordinary starlight photons at the point they enter the detector optics.
Since JWST is currently in space, such an adjustment is not feasible. However, ground-based tests of this kind could be implemented for similar future space telescopes such as LUVOIR [74], HabEx [37], Origins [60], Roman (WFIRST) [71], etc. In the next section, we will present more details on the ground test based on a modified JWST-like configuration.
Figure 2: The optical path inside the detector optics of NISpec onboard JWST. Figure taken from Ref. [51].
@@ 4 Projected Sensitivity with a Modified JWST Configuration
The arrangement of JWST’s mirrors has been optimized for observing distant astrophysical objects, making it not ideal for the purpose of observing the DPDM signal. This is the main reason that DPDM-induced photons do not contribute to the final detected signal, although such photons can indeed be generated within the telescope. Therefore, we propose to use the JWST primary mirror for a ground test. If we can adjust the relative positions of the mirrors, it would be possible to focus the DPDM-induced signal generated on the primary mirror onto the detector. As we have emphasized, such modification is only within a few mirrors in the OTE system while leaving the detector optics unchanged. Below, we analyze this possibility and the projected sensitivity limits under the modified mirror configuration. We will demonstrate that the state-of-the-art technology employed in JWST is exceptionally powerful for detecting novel or unconventional signals. Such predicted sensitivities can be potentially realized with the future space telescopes such as LUVOIR [74], HabEx [37], Origins [60], Roman (WFIRST) [71] during their pre-launch ground-testing phase. We use the parameters of JWST to illustrate the potential of such ground-based tests in searching for DPDM.
@@
Firstly, we simplify the OTE of JWST to an equivalent system of lenses. Without loss of generality, we assume that the direction of light rays is from left to right, while at the same time we replace each reflector with a corresponding type of lens for convenience in analysis, as illustrated in Figure 3. The radii of curvature for the first three lenses are denoted by $\rho_{1}$, $\rho_{2}$, and $\rho_{3}$, respectively. In this section, we will use subscripts $P$, $S$, $T$, and $F$ for the shorthand of “primary mirror”, “secondary mirror”, “tertiary mirror”, and “fine steering mirror”, respectively.
In the ray-transfer-matrix method, a ray is described with a 2-component vector $X$, the first component being the angle between the ray and the optical axis, the second being its vertical displacement from the optical axis. Each optical operation that a ray undergoes—such as free travel or refraction through a lens—is represented by a $2\times 2$ matrix. Specifically, within this section, free travel over a distance $L_{i}$ will be symbolized by the matrix $U_{i}$, while refraction on the primary mirror will be denoted by the matrix $U_{P}$, and likewise for other mirrors in the optical system.
@@
Light is emitted from each reflector. Since the primary mirror is the largest in size, we focus on the light generated on the primary mirror. The corresponding vector is
@@
Here, $y$ represents the height of the emission point relative to the optical axis.
@@
Since the light from distant stars is transformed into a thin collimated beam of light before being picked up by the detector, we should require the same for the DPDM signal. This way, we can ensure that the DPDM-generated signal in ground test would behave identically to the ordinary starlight in space missions, for which the detectors are designed. In the language of ray-transfer-matrix, we should require that after a series of reflections and propagation, the first component of the ray vector $U_{T}U_{2}U_{S}U_{1}X_{P}$ should be independent of $y$. This gives
@@
We also hope that the signal is not lost in the middle, which means all the light generated on the primary mirror must be picked up by the secondary mirror, tertiary mirror, and fine steering mirror. This gives us a set of equations,
@@
Where $D_{P}$, $D_{S}$, and $D_{F}$ are the diameters of the primary mirror, secondary mirror, and fine steering mirror, $A_{T}$ and $B_{T}$ are the length and width of the tertiary mirror. Combining Eq. (4.4), we find that these inequalities are satisfied with
@@
Figure 4: Constraints on the kinematic mixing parameter $\epsilon$ between Dark Photon Dark Matter (DPDM) and photons in the randomized polarization scheme. The dashed red curve shows the projected sensitivity for the modified JWST configuration. The left and right sections display constraints derived from NIRSpec and MIRI observation data, respectively. Additionally, we provide a comparison with existing limits, including those from Solar [54], XENON1T [13], Lampost [23], Mudhi [58], Funk [12], and Tokyo [72].
When all the signals generated on the whole primary mirror can be detected, it’s straightforward to determine the equivalent flux, which equals the total power of the signal divided by the area of the primary mirror. This means the equivalent flux is directly given by Eq. (2.24),
@@
This value, divided by the bandwidth, should be compared with the designed sensitivity of spectral flux density. The projection curves are shown as the red dashed curves in Fig. 4. Since the faintest signal in JWST data is from the observation of the cold sky, we directly use this data as an estimation of the future ground test background, assuming the ground test will be performed with a background temperature comparable to the cosmic radio background. A lower background temperature would give us a stronger constraint on $\epsilon$.
Figure 5: Model-independent 95% C.L. upper limits on a constant monochromatic signal from JWST data. The first figure corresponds to NIRSpec observation data, and the lower one to MIRI. It shows the strongest limit from all the projects at each frequency bin.
In our study, we harnessed data from 972 distinct observation projects to establish constraints on the Dark Photon-Dark Matter coupling constant. Out of these projects, 713 relied on the Near-Infrared Spectrograph (NIRSpec) [49], while 259 made use of the Mid-Infrared Instrument (MIRI) [68]. It is worth noting that the data selected for analysis excludes background subtraction, ensuring its suitability for our research.
The JWST data we collected from the Mikulski Archive for Space Telescopes (MAST) database [1] includes two crucial parameters: the measured spectral flux density, denoted as $\bar{O}_{i}$, and the associated statistical uncertainty, denoted as $\sigma_{\bar{O}_{i}}$. In our effort to establish upper limits on $\epsilon$, the coupling of dark photon with the Standard Model electromagnetic current, we followed the data analysis approach detailed in previous works [26, 9, 8].
To provide a concise overview, we summarize the key aspects of our data-analysis method here, while reserving more detailed information for Appendix D. Our analysis begins by applying a local polynomial function to model the background surrounding a selected frequency bin, $i_{0}$, while considering neighboring bins. We estimate systematic uncertainties by comparing data deviations to the background fit. Next, we introduce a hypothetical DPDM-induced signal with a strength denoted as $S$ at the specific bin, $i_{0}$. This allows us to construct a likelihood function, $L$, that incorporates $S$ into the comparison between data and the background function. Nuisance parameters are introduced to account for the coefficients of the background polynomial function.
Following the statistical method developed in Ref. [26], we compute the ratio, $\lambda_{S}$, between the maximized likelihood under two conditions: first, when only the nuisance parameters are varied to maximize $L$ while keeping $S$ constant, and second, when both the nuisance parameters and $S$ are varied to maximize $L$. The test statistic, $2\log(\lambda_{S})$, follows a half-$\chi^{2}$ distribution [26]. This analysis allows us to derive the $95\%$ C.L. upper limit, $S_{\mathrm{lim}}$, for a constant monochromatic signal. The results are illustrated in Fig. 5.
We establish upper limits on the mixing parameter $\epsilon$ as $S_{\mathrm{lim}}=S^{\prime}_{\mathrm{eqv}}$, where $S^{\prime}_{\mathrm{eqv}}$ represents the signal strength from theoretical calculations for DPDM. $S^{\prime}_{\mathrm{eqv}}=I^{\mathrm{proj}}_{\mathrm{eqv}}/\mathcal{B}$ where $I^{\mathrm{proj}}_{\mathrm{eqv}}$ is from (4.11) and $\mathcal{B}$ is the bandwidth of JWST. Different datasets from NIRSpec and MIRI yield varying constraints on the signal strength coupling $\epsilon$, and we select the most stringent among them. The predicted constraints are shown in Fig. 4 as red dashed curves, along with the existing constraints.
Indeed, the configuration of JWST, now operational in space, cannot be modified to optimize its sensitivity to DPDM. However, for similar future space telescopes (e.g., LUVOIR [74], HabEx [37], Origins [60], Roman (WFIRST) [71]), it may be possible to adjust the optical configuration during the ground-testing phase to enhance their suitability for DPDM detection. Then, these telescopes would be able to offer significant sensitivity to DPDM, as demonstrated in the projected sensitivity shown in Fig. 4 for the JWST case.
@@ 5 Summary and Outlook
In this study, we have conducted a direct detection search for DPDM using a haloscope setup. We explore direct detection of local DPDM through its conversion into a conventional electromagnetic field at JWST. Due to the frequency match between the electromagnetic field and the DPDM mass, the resulting signal took the form of a monochromatic electromagnetic wave. Moreover, we have proved that for high-mass DPDM, the propagation of the DPDM-induced signal follows simple ray optics, which greatly simplifies the calculation.
We have carefully analyzed the optical system of JWST, including both its OTE system and detector optics. Although DPDM can induce photons on multiple optical elements of JWST, these photons do not contribute to the final detected signal. The main reason is that DPDM-induced photons propagate nearly perpendicular to the local metallic surfaces, deviating from the optical path of starlight photons, and will be lost upon reflections. Nevertheless, we propose that, with appropriate adjustments to the mirror positioning in the OTE system during ground-based testing, a space telescope such as JWST would achieve significant sensitivity to DPDM. This concept can be applied to future space missions such as LUVOIR [74], HabEx [37], Origins [60], Roman (WFIRST) [71].
Using the parameters of NIRSpec and MIRI onboard JWST as representative examples, we calculate that the achievable upper limits on the DPDM–photon kinetic mixing parameter are $\epsilon\sim 10^{-12}-10^{-14}$ in the frequency range 1050010-500 THz at $95\%$ C.L. These results highlight the strong potential of future space telescopes to search for DPDM during their ground-testing phases by adopting the modified mirror configuration that enables DPDM-induced photons to be properly focused. Such a broadband search strategy would provide valuable constraints in the lower-frequency regime, complementing existing laboratory-based experiments such as Lampost, Mudhi, FUNK, and TOKYO. The expected sensitivity surpasses current limits by one to two orders of magnitude compared to the XENON1T result which utilized the potential solar dark-photon flux without assuming dark photons as dark matter, as well as the solar cooling bound.
Alternatively, during the ground-testing phase, it might be possible to employ a spherical mirror to generate the DPDM-induced photons and focus them before they enter the detector optics. If such a setup is feasible, the combined spherical mirror + detector optics system could achieve a sensitivity comparable to that of the modified OTE configuration, provided that the spherical mirror has an area similar to that of the primary mirror. Depending on which setup is more practical and experimentally accessible, one could choose either the modified OTE configuration or the spherical mirror + detector optics configuration for DPDM search experiments during ground testing.